OpenStudy (ffasinger):

A tangent-tangent angle intercepts two arcs that measure 135 and 225. What is the measure of the tangent-tangent angle? A. 180 B. 90 C. 45 D. 135

OpenStudy (jdoe0001):

got pic?

OpenStudy (jdoe0001):

either way -> http://www.mathwarehouse.com/geometry/circle/images/power-of-point/far-arc-near-arc-formula-picture-small.png check the "far arc near arc" equation

OpenStudy (ffasinger):

no pic

OpenStudy (janu16):

Angle x = (far arc - near arc) / 2 = (arc ACB - arc ADB) / 2 = (225 - 135) / 2 = 90 / 2 = 45 degrees.

OpenStudy (ffasinger):

thank you